Main INDEX
Monthly INDEX
PREV
NEXT
User name Humensky
Log entry time 05:15:22 on June28,2004
Entry number 127185
keyword=Hall C IA fdbk and asym width
Updated Hall C IA to 1.575 ~9k events into run 2491 to correct 268ppm asymmetry.
Tails on the Hall C charge asymmetry distribution have gotten a lot smaller recently, and the rms has dropped to ~1400ppm from 3500ppm. I checked the Hall C laser lock again and noticed that it now shows some noise on its phase value, similar to the Hall A laser lock phase noise. Earlier tonight, when the width was large on Hall C, its laser phase noise value was not changing (whereas Hall A laser phase noise was). This can be checked by opening up the "Tools" screen from an Accelerator Main Menu.